# Red Pill and Blue Pill: Controllable Website Fingerprinting Defense via Dynamic Backdoor Learning

### v1

@@ 6.3 Open-world Evaluation
This subsection evaluates the defense performance of CWFD alongside 6 other strategies against four different attack models in the open-world setting, as shown in Fig. 3. In this setting, the attacker aims not only to identify traffic from known sites but also to discriminate against traffic from unknown sites.
@@
Key findings from the analysis are as follows: ❶ CWFD consistently demonstrates superior defense capabilities across all types of attacks, particularly against the RF attack. CWFD (heavy) reduces the precision from a high of 0.9958 and recall from 0.9804 to nearly zero (0.0007 and 0.0005, respectively). CWFD (light) also effectively controls both precision and recall within a maximum range of 0.12. ❷ Comparison with other methods. CWFD (heavy) and CWFD (light) achieve lower mAP values compared to the state-of-the-art defenses Tamaraw and Palette in the open-world setting, showing significant advantages. Specifically, CWFD (heavy) achieves an mAP of 0.00025, while CWFD (light) achieves 0.0031, both outperforming Tamaraw (0.0060) and Palette (0.0061). Both CWFD (light) and CWFD (heavy) outperform ALPaCA (0.20), especially under RF attacks, CWFD achieves lower recall and precision. ❸ Broad and consistent defense. CWFD’s robust performance across multiple attack methods is notable. Unlike other strategies, which may perform well under certain conditions but falter in others, CWFD consistently maintains low precision (not exceeding 0.11) and recall (not exceeding 0.12) across all attack types. In contrast, RegulaTor, while performing adequately under the VarCNN attack, shows only mediocre results under RF and DF attacks, highlighting CWFD’s consistent reliability.
6.4 Real-World Performance
@@
CWFD 103.5 1.2 6.6 2.2 8.8 12.1
We prototype CWFD using WFDefProxy [13] to obtain precise measurements of its overhead and performance. WFDefProxy is a generic platform for deploying website fingerprint defenses within the Tor network, allowing each defense to function as a pluggable transport that proxies Tor traffic.
Deployment details. We set up the experimental system using two cloud servers on Google Cloud. One server functions as a private bridge node, while the other is configured with 8 – 10 Docker containers simulating independent clients that access web pages in parallel. The bridge server is equipped with a 2-core CPU and 4GB of RAM, while the client server has an 8-core CPU and 16GB of RAM. Both servers run on Ubuntu 20.04 LTS. The two servers are placed in different countries to create some physical distance. On each client, we run a Tor Browser with version 12.0.1. TBSelenium is used to automatically control the browser to visit specified web pages, with a maximum load time of 70 seconds for each visit. Each trace is collected over a different circuit to ensure that we capture the randomness of the Tor network.
TABLE III: The overhead and performance of CWFD under different network conditions.
@@
Real-world results. Tab. II shows the performance and overhead of CWFD compared to other defenses in a real-world setting. Notably, CWFD has a minimal time overhead (TO) of only 1.2%, significantly lower than methods like RegulaTor (58.4%) and Tamaraw (34.9%), making it more practical for real-time applications. In terms of data overhead (DO), CWFD maintains a moderate level of 63.4%, which is lower than FRONT (105.4%) and Tamaraw (113.7%), but higher than methods like Palette (59.3%). In terms of defense accuracy, CWFD achieves the best results across all attack types. For example, under the RF and DF attacks, CWFD reduces the attacker’s accuracy to 6.6% and 2.2%, respectively, outperforming other methods such as Surakav (44.5% on RF) and FRONT (43.2% on DF). Similarly, CWFD maintains low accuracy for TikTok and VarCNN attacks at 8.8% and 12.1%, respectively, significantly lower than Palette, which achieves 37.2% and 12.6% on these attacks. Overall, CWFD balances overhead and effectiveness effectively, offering robust defense with minimal time impact, making it highly suitable for real-world deployment.
Results on different bandwidth constraints. Tab. III evaluates CWFD’s performance and overhead across different network bandwidths (80, 120, and 160 Mbps). For overhead consistency, CWFD maintains stable data overhead (DO) around 101-104% and low time overhead (TO) across all bandwidths, with a slight increase in TO at 120 Mbps (2.8%). This demonstrates CWFD’s efficiency in diverse network conditions. Moreover, CWFD achieves stronger defense with increased bandwidth, especially for DF and VarCNN attacks. For instance, DF attack accuracy drops to 2.3% at 160 Mbps, while TikTok and VarCNN accuracies improve from 11.2% and 14.3% at 80 Mbps to 8.5% and 11.9% at 160 Mbps, respectively.
6.5 Defenses against Adaptive Attacks
TABLE IV: Performance of CWFD under different adaptive attack methods.
@@ 6.5 Defenses against Adaptive Attacks
CWFD-s CWFD-d CWFD-s CWFD-d CWFD-s CWFD-d
Undefended 6.1 1.3 3.2 1.5 2.3 9.4
RR 11.2 8.7 15.9 14.1 19.5 7.2
DT 6.0 2.3 3.7 2.1 6.4 5.2
CT 8.4 22.1 2.6 1.1 6.8 4.7
CF 2.3 0.3 8.6 28.5 8.8 21.3
@@
In this subsection, we examine adaptive attack strategies against CWFD from both data and model perspectives. At the data level, attackers might infer the mechanism by which CWFD poisons the training data and attempt to counteract it by randomly removing incoming cells or distinguishing between clean and poisoned traces during training. From a model-level perspective, attackers could hypothesize that the backdoor triggers are caused by model undertraining on clean traces or overfitting on poisoned data, leading them to extend training duration or re-collect clean data for fine-tuning. ❶ Random removal. In this approach, the attacker knows the total number of packets and bursts inserted by CWFD, such as 20,000 incoming packets over seven bursts. They may randomly select seven positions for each training trace and remove 2857 incoming cells at each point, while maintaining the same training parameters as if CWFD had not interfered. ❷ Discriminative training. If attackers suspect the presence of anomalous data, they may train a one-class SVM to differentiate between poisoned and clean data, using majority classification to filter out anomalies and keeping training parameters unchanged despite CWFD’s influence. ❸ Continuous training. Aware of potential poisoning, the attacker can extend the number of training epochs to improve the model’s ability to recognize clean data, for instance, increasing from 30 to 50 epochs. ❹ Clean fine-tuning. To counter CWFD’s influence and reduce reliance on poisoned data, an attacker could collect a small amount of clean data and fine-tune the backdoored model. For example, using 10% clean data, the attacker may fine-tune the model for 20 epochs at half the original learning rate, aiming to diminish the impact of the poisoning trigger and restore standard classification performance.
@@
Results and Analysis. Tab. IV shows the performance of CWFD under various adaptive attacks: Random Removal (RR), Discriminative Training (DT), Continuous Training (CT), and Clean Fine-tuning (CF). ❶ CWFD-d demonstrates strong resilience across adaptive attacks. CWFD-d maintains low attack accuracy in RF and DF scenarios, with only slight increases under RR and DT, indicating robustness against packet manipulation and discriminative tactics. ❷ CWFD-s provides better baseline defense but is more affected by CT and CF. While CWFD-s shows lower attack accuracy in the non-adaptive setting, it is more susceptible to adaptive methods, especially CF, which raises DF accuracy to 8.6%. ❸ Clean Fine-tuning is the most effective adaptive attack against CWFD-d. CF notably increases attack accuracy, particularly in DF (28.5%) and TikTok (21.3%), suggesting that access to clean data for fine-tuning can weaken CWFD-d’s defenses.
@@
Overall, CWFD-d’s adaptability provides superior resilience to most adaptive attacks, with clean fine-tuning as the main vulnerability.
6.6 Ablation Studies
@@
Figure 4: Effect of total incoming packets and burst count on CWFD defense effectiveness.
@@
Figure 5: Impact of poisoning rate (pr) on CWFD defense effectiveness across varying packet totals.
@@
We here investigate the impact of key CWFD parameters on defense performance in closed-world scenarios.
@@
Total incoming packets. We analyze defense performance across eleven configurations, ranging from 1,000 to 20,000 inserted packets per trace, with seven bursts per trace and a fixed poisoning rate of 1.0%. Fig. 4 (b) illustrates a positive trend in defense effectiveness with an increase in inserted packets. This improvement is due to the added complexity introduced by extra packets, which complicates the attacker’s task while maintaining minimal overhead relative to the total traffic. Defense performance peaks at 20,000 packets, which we select as the preferred configuration.
@@
Burst of incoming packets. This parameter denotes the number of insertion locations per trace. We examine burst settings from 1 to 9. As shown in Fig. 4 (a), increasing the number of bursts initially improves defense by adding more interference points, but the impact plateaus beyond seven bursts, suggesting diminishing returns. Consequently, seven bursts provide an optimal balance of defense strength and operational efficiency.
@@
Poisoning rate. Defined as the proportion of modified traces for a targeted tag, we evaluate poisoning rates from 0.1% to 1.0%. Fig. 5 illustrates that at lower packet totals (below 4,000), poisoning rate variations have limited impact on defense. However, with higher packet totals, increased poisoning rates significantly enhance defense, especially at 1.0%, where attack success rates are minimized. This trend is consistent across attack types, with RF attacks displaying a steady response and DF attacks showing more variability. Thus, we adopt a poisoning rate of 1.0%, combined with the optimal settings of 20,000 packets and seven bursts, to achieve robust defense across multiple attack scenarios while ensuring efficiency. Nevertheless, the poisoning rate remains very low, making our defense feasible for successfully poisoning the attack in real-world scenarios.
@@
Figure 6: Impact of target label choice on CWFD defense stability and effectiveness.
@@
Different target labels. Fig. 6 illustrates the effect of different target labels on CWFD defense under DF attack and RF attack in the Sirinam dataset. To explore the stability of CWFD defense over multiple epochs, we calculated the defense success rate from round 28 to round 32, a total of five epochs, and formed box-and-line. From Fig. 6, we conclude that ❶ Choosing different target labels can cause significant differences in defense performance. For example, in the RF attack, labels like 2 and 26 maintain stable accuracy around 6%, while in the DF attack, label 47 shows much wider fluctuations, ranging from 4% to 7%. This variation suggests that certain labels result in more consistent defense performance, especially under RF attacks, while DF attacks are more sensitive to label choice. ❷ Stability of the defense on the RF attack is better than that of the DF attack. For instance, the accuracy distribution for RF attack labels remains tightly clustered around the mean, with minimal variance across labels. In contrast, the DF attack shows greater variability, as seen with labels like 11 and 33, where the accuracy fluctuates significantly. This indicates that RF defenses maintain more consistent performance across different labels, while DF defenses are more prone to instability.

### v2

@@ VI-C Open-world Evaluation
This subsection evaluates the defense performance of TrapFlow alongside 6 other strategies against four different attack models in the open-world setting, as shown in Fig. 3. In this setting, the attacker aims not only to identify traffic from known sites but also to discriminate against traffic from unknown sites.
@@
Key findings from the analysis are as follows: ❶ TrapFlow consistently demonstrates superior defense capabilities across all types of attacks, particularly against the RF attack. TrapFlow (heavy) reduces the precision from a high of 0.9958 and recall from 0.9804 to nearly zero (0.0007 and 0.0005, respectively). TrapFlow (light) also effectively controls both precision and recall within a maximum range of 0.12. ❷ Comparison with other methods. TrapFlow (heavy) and TrapFlow (light) achieve lower mAP values compared to the state-of-the-art defenses Tamaraw and Palette in the open-world setting, showing significant advantages. Specifically, TrapFlow (heavy) achieves an mAP of 0.00025, while TrapFlow (light) achieves 0.0031, both outperforming Tamaraw (0.0060) and Palette (0.0061). Both TrapFlow (light) and TrapFlow (heavy) outperform ALPaCA (0.20), especially under RF attacks, TrapFlow achieves lower recall and precision. ❸ Broad and consistent defense. TrapFlow’s robust performance across multiple attack methods is notable. Unlike other strategies, which may perform well under certain conditions but falter in others, TrapFlow consistently maintains low precision (not exceeding 0.11) and recall (not exceeding 0.12) across all attack types. In contrast, RegulaTor, while performing adequately under the VarCNN attack, shows only mediocre results under RF and DF attacks, highlighting TrapFlow’s consistent reliability.
VI-D Real-world Performance
@@
TrapFlow 103.5 1.2 6.6 2.2 8.8 12.1
We prototype TrapFlow using WFDefProxy [18] to obtain precise measurements of its overhead and performance. WFDefProxy is a generic platform for deploying website fingerprint defenses within the Tor network, allowing each defense to function as a pluggable transport that proxies Tor traffic.
Deployment details. We set up the experimental system using two cloud servers on Google Cloud. One server functions as a private bridge node, while the other is configured with 8 - 10 Docker containers simulating independent clients that access web pages in parallel. The bridge server is equipped with a 2-core CPU and 4GB of RAM, while the client server has an 8-core CPU and 16GB of RAM. Both servers run on Ubuntu 20.04 LTS. The two servers are placed in different countries to create some physical distance. On each client, we run a Tor Browser with version 12.0.1. TBSelenium is used to automatically control the browser to visit specified web pages, with a maximum load time of 70 seconds for each visit. Each trace is collected over a different circuit to ensure that we capture the randomness of the Tor network.
TABLE III: The overhead and performance of TrapFlow under different network conditions.
@@
From an ethical standpoint, TrapFlow does not modify or delay any real user packets, nor does it introduce user identifiers or user-specific signatures. The defense operates solely by injecting dummy packets, which are structurally indistinguishable from standard Tor padding and are immediately discarded at the receiving end, ensuring that user data and application semantics remain unaffected. TrapFlow is implemented on top of WFDefProxy and follows Tor’s standard pluggable transport interface, operating within the existing Tor design framework. Moreover, the introduced overhead is intentionally kept minimal, resulting in negligible user-perceived impact. As such, TrapFlow does not raise ethical concerns beyond those already considered for existing padding-based Tor defenses.
Real-world results. Tab. II shows the performance and overhead of TrapFlow compared to other defenses in a real-world setting. Notably, TrapFlow has a minimal time overhead (TO) of only 1.2%, significantly lower than methods like RegulaTor (58.4%) and Tamaraw (34.9%), making it more practical for real-time applications. In terms of data overhead (DO), TrapFlow maintains a moderate level of 63.4%, which is lower than FRONT (105.4%) and Tamaraw (113.7%), but higher than methods like Palette (59.3%). In terms of defense accuracy, TrapFlow achieves the best results across all attack types. For example, under the RF and DF attacks, TrapFlow reduces the attacker’s accuracy to 6.6% and 2.2%, respectively, outperforming other methods such as Surakav (44.5% on RF) and FRONT (43.2% on DF). Similarly, TrapFlow maintains low accuracy for TikTok and VarCNN attacks at 8.8% and 12.1%, respectively, significantly lower than Palette, which achieves 37.2% and 12.6% on these attacks. Overall, TrapFlow balances overhead and effectiveness effectively, offering robust defense with minimal time impact, making it highly suitable for real-world deployment.
Results on different bandwidth constraints. Tab. III evaluates TrapFlow’s performance and overhead across different network bandwidths (80, 120, and 160 Mbps). For overhead consistency, TrapFlow maintains stable data overhead (DO) around 101-104% and low time overhead (TO) across all bandwidths, with a slight increase in TO at 120 Mbps (2.8%). This demonstrates TrapFlow’s efficiency in diverse network conditions. Moreover, TrapFlow achieves stronger defense with increased bandwidth, especially for DF and VarCNN attacks. For instance, DF attack accuracy drops to 2.3% at 160 Mbps, while TikTok and VarCNN accuracies improve from 11.2% and 14.3% at 80 Mbps to 8.5% and 11.9% at 160 Mbps, respectively.
System tradeoffs regarding DO vs. TO. The slightly higher overhead observed in the real implementation of TrapFlow compared to simulation results is expected and consistent with prior work. This difference mainly arises because real systems lack access to precise traffic boundary information (e.g., exact session start and end times) and must instead infer session phases using finite-state machines, which inevitably introduces limited redundant trigger injection. Similar implementation–simulation gaps have also been reported in prior studies [18, 61]. From a system perspective, time overhead rather than data overhead is the primary constraint in Tor deployments, as even small delays can significantly degrade user experience and network stability. Accordingly, TrapFlow deliberately trades a modest increase in data overhead for near-zero time overhead. As shown in Tab. II, although TrapFlow incurs higher DO than Palette, its TO is essentially negligible while providing substantially stronger protection against all evaluated attacks, representing a more favorable system-level tradeoff in real Tor environments.
VI-E Defenses against Adaptive Attacks
TABLE IV: Performance of TrapFlow under different adaptive attack methods.
@@ VI-E Defenses against Adaptive Attacks
s d i s d i s d i
Undefended 4.7 5.8 6.2 2.8 3.9 4.1 3.4 4.7 5.3
$R$ 11.2 8.7 8.5 15.9 14.1 14.3 19.5 7.2 7.2
DT 6.0 2.3 2.1 3.7 2.1 2.0 6.4 5.2 4.8
CT 2.3 0.3 0.3 2.6 1.1 0.9 6.8 4.7 4.6
CF 8.4 22.1 9.2 8.6 28.5 9.6 8.8 21.3 9.7
@@
In this subsection, we examine adaptive attack strategies against TrapFlow from both data and model perspectives. At the data level, attackers might infer the mechanism by which TrapFlow poisons the training data and attempt to counteract it by randomly removing incoming cells or distinguishing between clean and poisoned traces during training. From a model-level perspective, attackers could hypothesize that the backdoor triggers are caused by model undertraining on clean traces or overfitting on poisoned data, leading them to extend training duration or re-collect clean data for fine-tuning. We elaborate on the design principles in Supplementary Materials G. ❶ Random removal. In this approach, the attacker knows the total number of cells and bursts inserted by TrapFlow, such as 20,000 incoming cells over seven bursts. They may randomly select seven positions for each training trace and remove 2857 incoming cells at each point, while maintaining the same training parameters as if TrapFlow had not interfered. ❷ Discriminative training. If attackers suspect the presence of anomalous data, they may train a one-class SVM to differentiate between poisoned and clean data, using majority classification to filter out anomalies and keeping training parameters unchanged despite TrapFlow’s influence. ❸ Continuous training. Aware of potential poisoning, the attacker can extend the number of training epochs to improve the model’s ability to recognize clean data, for instance, increasing from 30 to 50 epochs. ❹ Clean fine-tuning. To counter TrapFlow’s influence and reduce reliance on poisoned data, an attacker could collect a small amount of clean data and fine-tune the backdoored model. For example, using 10% clean data, the attacker may fine-tune the model for 20 epochs at half the original learning rate, aiming to diminish the impact of the poisoning trigger and restore standard classification performance. Although our adaptive attack strategies are proposed from the attacker’s perspective, they inherently mirror several ideas from classical backdoor defense literature.
@@
Results and analysis. Tab. IV shows the performance of TrapFlow under various adaptive attacks: Random Removal (RR), Discriminative Training (DT), Continuous Training (CT), and Clean Fine-tuning (CF). ❶ TrapFlow-d demonstrates strong resilience across adaptive attacks. TrapFlow-d maintains low attack accuracy in RF and DF scenarios, with only slight increases under RR and DT, indicating robustness against packet manipulation and discriminative tactics. ❷ TrapFlow-s provides better baseline defense but is more affected by CT and CF. While TrapFlow-s shows lower attack accuracy in the non-adaptive setting, it is more susceptible to adaptive methods, especially CF, which raises DF accuracy to 8.6%. ❸ Clean Fine-tuning is the most effective adaptive attack against TrapFlow-d. CF notably increases attack accuracy, particularly in DF (28.5%) and TikTok (21.3%), suggesting that access to clean data for fine-tuning can weaken TrapFlow-d’s defenses.
@@
Overall, TrapFlow-d’s adaptability provides superior resilience to most adaptive attacks, with clean fine-tuning as the main vulnerability. Although our adaptive attack strategies are proposed from the attacker’s perspective, they inherently mirror several ideas from classical backdoor defense literature.
CF-Improved trigger design. We view CF as an incremental adapt-and-forget process with a limited budget. Specifically, using a small amount of clean data, the attacker locally corrects the model that has been affected by the defense perturbation in an attempt to forget the bias introduced by the defense. Whether CF is successful critically depends on whether the bias can be selectively removed without destroying the task discriminative structure; when the bias is manifested only as a weakly correlated noise feature of the task, clean fine-tuning tends to suppress it efficiently, whereas when the bias is intertwined with the discriminative statistical structure of the target category, forcibly forgetting the bias will inevitably compromise the discriminative power that the attacker aims to recover, leading to incomplete forgetting. This perspective is consistent with the motivation of BadCLIP [37], which resists subsequent corrections by enhancing the coupling between trigger signals and target semantic features.
Based on this insight, we propose an improved trigger design(denoted as i in Tab. IV). While maintaining defense perturbations with sufficient edit differences (constrained by the fast Levenshtein distance $D_{F}$) in the input sequence space to ensure defense controllability, we further introduce target-label statistical alignment, such that the defended sequence retains a portion of the target-label features in terms of key statistical structure. Specifically, let the original traffic sequence be $\bm{x}$, the post-defense sequence be $\hat{\bm{x}}$, and the target label be $\eta(y)$. The CF-improved trigger can then be formulated as
$\bm{\delta}^{\text{CF}}=\arg\min_{\delta\in\Omega}\big|S(\hat{\bm{x}})-\mu_{\eta(y^{\star})}\big|_{2}\quad\text{s.t.}\quad D_{F}(\bm{x},\hat{\bm{x}})\geq\tau,$ (9)
where $S(\cdot)$ denotes the burst-level statistical features (e.g., low-order statistics of burst length and direction distributions) extracted from the traffic sequence, and $\mu_{\eta(y^{\star})}$ represents the statistical prototype of the target label computed from clean data.
Tab. IV systematically compares the defense effectiveness of TrapFlow under multiple adaptive attack strategies, where s denotes TrapFlow-s (heavy), d denotes TrapFlow-d (heavy), and i denotes the improved version of TrapFlow-d that incorporates the CF-improved trigger. Overall, all three defense configurations maintain attack accuracy at a low level under adaptive attacks such as Random Removal (RR), Discriminative Training (DT), and Continuous Training (CT), indicating that TrapFlow is robust to common adaptive strategies. Among them, the i-version performs comparably to the d-version in most settings, with slight improvements in some scenarios, indicating that the introduction of statistical alignment constraints does not weaken the overall defense capability and does not sacrifice the original defense strength.
The differences among the configurations become particularly pronounced under CF attacks. For TrapFlow-d, CF significantly degrades the defense effectiveness, e.g., the attack accuracy rises to 28.5% under the DF attack and reaches 22.1% and 21.3% under the RF and TikTok attacks, respectively. In contrast, the i-version, which introduces the CF-improved trigger, effectively suppresses this performance regression: the accuracy is maintained at 9.6%, 9.2%, and 9.7% under the DF, RF, and TikTok attacks, respectively, which is substantially lower than that of the d-version and remains at a level comparable to that of the s-version.
These results indicate that the CF-improved trigger more tightly couples the trigger perturbation with the discriminative statistical structure of the target category by introducing alignment with target-label statistics while preserving sufficient edit differences (constrained by $D_{F}$). When the attacker attempts to “forget” the defense perturbation via clean fine-tuning, the induced bias no longer behaves as an independently removable noise feature, but instead becomes entangled with the discriminative power of the target category, thereby substantially reducing the effectiveness of the CF attack.
VI-F Ablation Studies
@@
Figure 4: Effect of total incoming cells and burst count on TrapFlow defense effectiveness.
@@
Figure 5: Impact of poisoning rate (pr) on TrapFlow defense effectiveness across varying packet totals.
@@
We here investigate the impact of key TrapFlow parameters on defense performance in closed-world scenarios.
@@
Total incoming cells. We analyze defense performance across eleven configurations, ranging from 1,000 to 20,000 inserted cells per trace, with seven bursts per trace and a fixed poisoning rate of 1.0%. Fig. 4 (b) illustrates a positive trend in defense effectiveness with an increase in inserted cells. This improvement is due to the added complexity introduced by extra cells, which complicates the attacker’s task while maintaining minimal overhead relative to the total traffic. Defense performance peaks at 20,000 cells, which we select as the preferred configuration.
@@
Burst of incoming cells. This parameter denotes the number of insertion locations per trace. We examine burst settings from 1 to 9. As shown in Fig. 4 (a), increasing the number of bursts initially improves defense by adding more interference points, but the impact plateaus beyond seven bursts, suggesting diminishing returns. Consequently, seven bursts provide an optimal balance of defense strength and operational efficiency.
@@
Poisoning rate. Defined as the proportion of modified traces for a targeted tag, we evaluate poisoning rates from 0.1% to 1.0%. Fig. 5 illustrates that at lower packet totals (below 4,000), poisoning rate variations have limited impact on defense. However, with higher packet totals, increased poisoning rates significantly enhance defense, especially at 1.0%, where attack success rates are minimized. This trend is consistent across attack types, with RF attacks displaying a steady response and DF attacks showing more variability. Thus, we adopt a poisoning rate of 1.0%, combined with the optimal settings of 20,000 cells and seven bursts, to achieve robust defense across multiple attack scenarios while ensuring efficiency. Nevertheless, the poisoning rate remains very low, making our defense feasible for successfully poisoning the attack in real-world scenarios.
@@
Figure 6: Impact of target label choice on TrapFlow defense stability and effectiveness.
@@
Different target labels. Fig. 6 illustrates the effect of different target labels on TrapFlow defense under DF attack and RF attack in the Sirinam dataset. To explore the stability of TrapFlow defense over multiple epochs, we calculated the defense success rate from round 28 to round 32, a total of five epochs, and formed box-and-line. From Fig. 6, we conclude that ❶ Choosing different target labels can cause significant differences in defense performance. For example, in the RF attack, labels like 2 and 26 maintain stable accuracy around 6%, while in the DF attack, label 47 shows much wider fluctuations, ranging from 4% to 7%. This variation suggests that certain labels result in more consistent defense performance, especially under RF attacks, while DF attacks are more sensitive to label choice. ❷ Stability of the defense on the RF attack is better than that of the DF attack. For instance, the accuracy distribution for RF attack labels remains tightly clustered around the mean, with minimal variance across labels. In contrast, the DF attack shows greater variability, as seen with labels like 11 and 33, where the accuracy fluctuates significantly. This indicates that RF defenses maintain more consistent performance across different labels, while DF defenses are more prone to instability. In addition, please see Supplementary Materials I for details on the performance of the defense algorithm on the DS-19 dataset. We also present the visualization results of TrapFlow in Supplementary Material J.
TABLE V: Performance comparison under DF and RF attacks with and without agnostic setting.
Attack Overhead (%) $\downarrow$ Accuracy (%) $\downarrow$
DO TO 30 50 all
DF 73.7 0.0 $4.0\pm 0.73$ $4.1\pm 0.72$ $4.0\pm 0.75$
RF 73.7 0.0 $5.9\pm 0.46$ $6.1\pm 0.48$ $6.0\pm 0.51$
DF (agnostic) 85.8 0.0 $4.1\pm 0.42$ $4.1\pm 0.40$ $4.1\pm 0.40$
RF (agnostic) 85.8 0.0 $6.0\pm 0.26$ $6.0\pm 0.25$ $5.0\pm 0.26$
Label agnostic stabilization. In the previous paragraph, we observed that the TrapFlow defense exhibits noticeable variations across different target labels, and that these variations are particularly pronounced under DF attacks. Further analysis reveals that this target-label dependence is closely related to differences in traffic length across websites. Specifically, for websites with shorter traffic sequences, the relative proportion of triggering patterns within the overall sequence is higher, making the backdoor features more likely to be learned in a stable manner and thus resulting in stronger defense effects; whereas for websites with longer traffic sequences, the same triggering patterns are more easily diluted by the original traffic structure, leading to weaker defense effects and larger performance fluctuations.
To address the above issue, we propose a label-agnostic defense stabilization strategy. Instead of relying on the selection of specific target labels, we design the trigger pattern in a uniform manner based on the statistical characteristics of traffic. Specifically, we adopt a dynamic injection mechanism normalized by traffic length, inserting the trigger pattern at an average interval of approximately 2500 packets and introducing a random jitter of $\pm 300$ packets around each insertion position, so as to ensure that web traffic of different lengths is subjected to approximately the same relative perturbation strength. This defense variant is denoted as agnostic in Tab. V.
The experimental results are shown in Tab. V. We report the defense effectiveness under DF and RF attacks on the Sirinam dataset, where 30, 50, and all denote the cases of randomly selecting 30 target labels, randomly selecting 50 target labels, and evaluating all target labels, respectively. As can be seen, the proposed agnostic strategy significantly reduces performance fluctuations across different target labels while keeping the average defense effectiveness largely unchanged. Specifically, under the DF attack, the cross-label standard deviation is reduced from approximately 0.72-0.75 under the baseline heavy setting to about 0.40, indicating a substantial improvement in defense stability; under the RF attack, the standard deviation is similarly reduced from about 0.46-0.51 to about 0.25, which further validates the effectiveness of the strategy in mitigating target-label dependence. These results demonstrate that the insensitivity of the TrapFlow defense to target label selection is effectively enhanced by the length-normalized dynamic injection design.
VI-G Measuring Adversarial Misclassification Effect